OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = one thirdx + 7 y = −3x + 3 y = one thirdx + 3 y = −3x + 7

OpenStudy (anonymous):

@gypsy1274 I know I have to use the same, y-y1 =m (x-x1) process

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

The only difference between this one and the last one is the slope. Perpendicular requires a negative reciprocal. Parallel has the same slope.

OpenStudy (anonymous):

Yes, you are right.

OpenStudy (anonymous):

So this one is still Y- (-1) = 1/3 (x-6)?

OpenStudy (anonymous):

(x,y) (-1,6) Do you see your mistake?

OpenStudy (anonymous):

And this question is asking for a parallel line so the slope will be the same as the given line.

OpenStudy (anonymous):

Aw poop. I do, I am sorry.

OpenStudy (anonymous):

No problem. If at first you don't succeed, try, try again.

OpenStudy (anonymous):

Y-6= 1/3 (x - -1)?

OpenStudy (anonymous):

Every mistake provides an opportunity to learn.

OpenStudy (anonymous):

Almost. What is the slope of the original equation?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

y=mx+b m=slope

OpenStudy (anonymous):

y=-3x+12

OpenStudy (anonymous):

So I solve from there? Or...?

OpenStudy (anonymous):

Find the correct slope.

OpenStudy (anonymous):

What is the slope of the original line?

OpenStudy (anonymous):

Wait! I think I seen this problem in an example from the lesson!

OpenStudy (anonymous):

I am going to go out on a limb here and say it is Y= -3X =3?

OpenStudy (anonymous):

Do you understand how to do the problem?

OpenStudy (anonymous):

Getting the correct answer without understanding cheats you of learning.

OpenStudy (anonymous):

That was my photographic memory speaking. But not really, so I would really appreciate if you explained.

OpenStudy (anonymous):

Start by finding the slope of the original line. y=-3x+12 y=mx+b m=slope

OpenStudy (anonymous):

I am stumped on how to find the slope, what numbers do I include, and how do I get them?

OpenStudy (anonymous):

You do not need to do anything to it, just find it. Re-read the previous post and see if you can figure it out.

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

So can I add1? seeing that is parallel?

OpenStudy (anonymous):

Have you identified the slope of the original equation?

OpenStudy (anonymous):

No, not really, slope always gets me. I know the way I need to set it up, but I never know HOW someone gets it. What do they do to get that number.

OpenStudy (anonymous):

You don't do anything to it. Just find it. m=slope. m is the coefficient of the x term.

OpenStudy (anonymous):

Is it making sense yet?

OpenStudy (anonymous):

So the slope is 12, or -3?

OpenStudy (anonymous):

I am thinking -3.

OpenStudy (anonymous):

The slope is -3.

OpenStudy (anonymous):

Ok, so we set it up like y-6 = -3 (x- -1)

OpenStudy (anonymous):

A parallel line has the same slope. So plug the point values and the slope in to the point-slope formula.

OpenStudy (anonymous):

Ok, so we set it up like y-6 = -3 (x- -1)

OpenStudy (anonymous):

Yes. Then solve for y.

OpenStudy (anonymous):

Ok, so when we solve for Y, we distribute -3?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

which leaves us with y- 6= -3x -3?

OpenStudy (anonymous):

yes. One more step.

OpenStudy (anonymous):

We get rid of six? By subtracting on both sides?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So then we have y= -3x + 3!

OpenStudy (anonymous):

Yes. You got it! \[\Huge \text{(^◡^ )}\]

OpenStudy (anonymous):

Yay! ok, I am going to try the next problem myself, then post it here for you to check, if I get it right, then Woo! If not ,then you can explain :D?

OpenStudy (anonymous):

Sounds like a great plan.